Can someone help me on this problem?
I need to solve the following equation for x, and show proof.
2(x-3) -3(x+5) = 3(x-2) -7
Thanks,
Sher

Answer by checkley71(8403)

2(x-3)-3(x+5)=3(x-2)-7 or 2x-6-3x-15=3x-6-7 or 2x-3x-3x=-6-7+6+15 or
-4x=8 or x=8/-4 or x=-2
proof 2(-2-3)-3(-2+5)=3(-2-2)-7 or 2(-5)-3(3)=3(-4)-7 or -10-9=-12-7 or -19=-19